OpenStudy (anonymous):

Can some help me out with this ?? porfavorcitow Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11

OpenStudy (anonymous):

Have you learnt matrices yet?

OpenStudy (anonymous):

and echelon form?

OpenStudy (anonymous):

yes I already did everything but dont know how to explain it step by step

OpenStudy (anonymous):

x - 5y + 2z = 0 x=5y - 2z 5y - 2z + 4y - z = 12 9y-3z=12 -3z=12-9y z=(12-9y)/-3 z=-4+3y 2x – y + 3z = 10 2(5y - 2z) - y + 3z = 10 10y - 4z - y +3z = 10 9y - z= 10 9y - (-4+3y) = 10 9y+4-3y=10 6y=10-4 6y=6 y=6/6 y=1 z=-4+3(1) z=-4+3 z=-1 x + 4(1) – (-1) = 12 x+4+1=12 x+5=12 x=12-5 x=7 7 - 5(1) + 2(-1) = 0 7-5-2=0 7-7=0 7+ 4(1) – (-1) = 12 7+4+1=12 7+5 = 12 2(7) – 1 + 3(-1) = 10 14-1-3=10 14-4=10

OpenStudy (anonymous):

hmm. I don't quite know how you solved for it.  that's what i'd have done. and the explanation for that is that you can use just the coefficients to solve for x y and z since they correlate to each other.